OpenStudy (anonymous):

Look at this equation. w = 2 /3t + 6 What is the value of w when t = 12 ? . A 24 B 14 C 12 D 9

OpenStudy (anonymous):

B

OpenStudy (texaschic101):

is it (2/3)t or 2/(3t) ?

OpenStudy (anonymous):

Just plug in the value given for t and solve...

OpenStudy (texaschic101):

lets do both... w = (2/3)t + 6 w = 2/3(12) + 6 w = 8 + 6 w = 14 or w = 2/(3t + 6) w = 2/(3(12) + 6) w = 2/(36 + 6) w = 2/42 w = 1/24 nvm....I guess it was the first way. answer is B. 14

OpenStudy (anonymous):

thanx